# Geometric Time-Dependent Density Functional Theory

### v1

@@ Abstract
We provide a new formulation of Time-Dependent Density Functional Theory (TDDFT) based on the geometric structure of the set of states constrained to have a fixed density. Orbital-free TDDFT is formulated using a hydrodynamics equation involving a new universal density-to-current functional map. In the corresponding Kohn–Sham equation, the density is reproduced using a non-local operator. Numerical simulations for one-dimensional soft-Coulomb systems are provided.
Density Functional Theory (DFT) is very successful in describing electrons at equilibrium, that is, close to their ground states. It is nowadays the reference computational method for large molecules and solids, both in academic research and industrial applications. By contrast, Time-Dependent Density Functional Theory (TDDFT) is less developed and still suffers from important limitations [1, 2, 3]. It is however the only viable computational approach for complex systems at present, and the demand for novel concepts and approximations in TDDFT is therefore considerable. Practical applications of TDDFT include charge transfer in complex molecules [4], electronic response of large systems [5], and ultrafast phenomena in attosecond physics [6]. In this Letter we open a new avenue in TDDFT. Based on geometric concepts, we completely revisit its theoretical foundations. We obtain a new TDDFT formulation that seems to better describe systems driven out of equilibrium. We also present 1D numerical simulations that illustrate the potential power of our new approach, when used as a correction to adiabatic density functionals.
@@
where $\widehat{T}:=-\sum_{j=1}^{N}\nabla^{2}_{\mathbf{r}_{j}}/2$ is the kinetic-energy operator, $\widehat{V}_{\rm ee}$ is the electron-electron interaction and $\widehat{V}_{\rm ext}(t):=\sum_{j=1}^{N}V_{\rm ext}(t,\mathbf{r}_{j})$ is a given time-dependent external potential. The reference $N$-particle Schrödinger equation is
@@
The goal of orbital-free TDDFT is to replace the high-dimensional linear equation (2) by an equivalent nonlinear equation involving only the one-particle density
@@
To this end, we introduce a new approach that consists in embedding Eq. (2) into a family of modified Schrödinger equations, parametrized by the initial state $\Psi_{0}$ and a prescribed time-dependent density $\rho(t,\mathbf{r})$ starting at $\rho(0,\mathbf{r})=\rho_{\Psi_{0}}(\mathbf{r})$. Those take the general form
$i\partial_{t}\Psi(t)=\Big(\widehat{H}(t)+\widehat{F}[\Psi_{0},\rho](t)\Big)\Psi(t),\quad\Psi(0)=\Psi_{0},$ (4)
@@
with a correction term $\widehat{F}[\Psi_{0},\rho](t)$ to be determined such that $\rho_{\Psi(t)}(\mathbf{r})=\rho(t,\mathbf{r})$. To ensure causality, $\widehat{F}[\Psi_{0},\rho](t)$ is assumed to only depend on the past trajectory $\{\rho(s,\mathbf{r})\}_{0\leqslant s\leqslant t}$. There are many possible choices for the operator-valued functional $\widehat{F}$. Based on geometric considerations, we will construct in this Letter a functional for which the density $\rho^{\rm S}:=\rho_{\Psi^{\rm S}}$ of the Schrödinger wavefunction $\Psi^{\rm S}$ in (2) is the only solution to the implicit equation $\widehat{F}[\Psi_{0},\rho]=0$. This is how (2) can be reformulated in terms of the density only.
@@
Standard TDDFT can be reinterpreted as the choice $\widehat{F}^{\rm VP}[\Psi_{0},\rho](t):=\sum_{j=1}^{N}V(t,\mathbf{r}_{j})$, where the potential $V=V[\Psi_{0},\rho]$ is seen as a functional of $\Psi_{0}$ and $\rho$. This comes from requiring the action to be stationary [7, 8] and is often called the Time-Dependent Variational Principle (VP). The Runge–Gross theorem [9] states that $V$ is unique modulo a time-dependent constant, hence $\rho^{\rm S}$ is the unique solution to $\widehat{F}^{\rm VP}[\Psi_{0},\rho]=0$.
In our new Geometric Principle (GP) we do not demand the action to be stationary, but instead require that the norm of the correction, $\|\widehat{F}^{\rm GP}[\Psi_{0},\rho]\Psi(t)\|=\|i\partial_{t}\Psi(t)-\widehat{H}(t)\Psi(t)\|$, be the smallest it can be at all times. As we explain shortly, this leads to the modified Schrödinger equation
$i\partial_{t}\Psi(t)=\bigg(\widehat{H}(t)+i\sum_{j=1}^{N}W(t,\mathbf{r}_{j})\bigg)\Psi(t),$ (5)
with $W=W[\Psi_{0},\rho]$ a real functional of $\Psi_{0}$ and $\rho$. In this Letter we provide a new formulation of TDDFT based on (5) and discuss important consequences.
@@
We first briefly explain why (5) minimizes the norm of the correction and refer to [10] for details. For simplicity, we consider a prescribed time-independent density $\rho(\mathbf{r})$. The wavefunction $\Psi(t)$ in (4) then evolves on the set $\mathcal{M}_{\rho}$ of wavefunctions with this prescribed $\rho$ and the velocity $\partial_{t}\Psi(t)$ must belong to the tangent space $\mathcal{T}_{\Psi(t)}\mathcal{M}_{\rho}$ at all times. As shown in Fig. 1, the correction term has minimal norm when $\partial_{t}\Psi(t)$ is the projection of $i\widehat{H}(t)\Psi(t)$ onto the tangent space, i.e. when $i\widehat{F}^{\rm GP}[\Psi_{0},\rho]\Psi(t)$ is in the normal space $\mathcal{N}_{\Psi(t)}:=(\mathcal{T}_{\Psi(t)}\mathcal{M}_{\rho})^{\perp}$. By simple calculations [10], the normal space is
@@
and this is how to arrive at (5). The correction is
$\widehat{F}^{\rm GP}[\Psi_{0},\rho]\Psi(t):=i\widehat{W}(t)\Psi(t),\quad\widehat{W}(t):=\sum_{j=1}^{N}W(t,\mathbf{r}_{j}).$ (7)
For a prescribed time-dependent density, one should take into account the displacement of the tangent space in the normal direction [10], which does not change the final equation (5). The idea of constraining a Schrödinger dynamics using a projection on the tangent space, often attributed to McLachlan [11], can be traced back to Dirac [12] and Frenkel [13]. Depending on the considered constraints, the GP can lead to the same solution as the VP, or to completely different solutions [14, 15, 16, 17, 18, 19]. In TDDFT, the GP and VP are very different.
Equation (5) looks like a Schrödinger equation with a complex potential, similar to those considered in TDDFT for describing resonances [20, 21, 22, 23, 24] or open systems [25, 26, 27]. However, here, $W$ is chosen so that $\rho_{\Psi(t)}=\rho(t)$ and thus the resulting $\Psi(t)$ stays normalized for all times. To make (4) a nonlinear Schrödinger equation with a Hermitian Hamiltonian, the correction term appearing in (7) can in fact be defined as the commutator
$\widehat{F}^{\rm GP}[\Psi_{0},\rho](t):=i\Big[\widehat{W}(t)\,,\,|\Psi(t)\rangle\langle\Psi(t)|\Big],$ (8)
because $\langle\Psi(t),\widehat{W}(t)\Psi(t)\rangle=0$. In [10] we give a rigorous proof of the existence and uniqueness of $W[\Psi_{0},\rho]$ in finite dimension. In Appendix A we rigorously prove a Runge–Gross theorem for Eq. (5) with smooth potentials. Hence $W[\Psi_{0},\rho]$ is uniquely defined, when it exists, and $\rho^{\rm S}$ is the unique solution to $\widehat{F}^{\rm GP}[\Psi_{0},\rho]=0$, or equivalently $W[\Psi_{0},\rho]=0$.
The function $W$ is very different in nature from the potential $V$ used in standard TDDFT. In particular, the potential $V$ often contains high peaks and steps to force the particles to move to the desired region of space [28]. This only works thanks to a complicated interplay between $V$ and the kinetic-energy operator $\widehat{T}$. Indeed, if we remove $\widehat{T}$, then adding a real-valued local potential $V$ modifies the phase of the wavefunction but not the density. In constrast, the function $W$ appears directly in the continuity equation associated with (5)
@@
where $\mathbf{j}_{\Psi}$ and $\rho_{\Psi}^{(2)}$ are the current and the two-particle density of $\Psi$. The linear operator $\mathcal{L}_{\Psi}$ is expected to be invertible in all cases of physical interest [10]. The interpretation of (9) is that $W$ gives rise to sources and sinks [21], which is a much more direct way of changing the density. On a lattice, densities varying too fast are not $V$-representable due to the finite speed of propagation [22] but can perfectly be $W$-representable [10].
Let us now re-interpret the implicit equation $W[\Psi_{0},\rho]=0$ in light of the continuity equation (9). From the functional $W[\Psi_{0},\rho]$ we obtain a unique wavefunction $\Psi^{\rm GP}[\Psi_{0},\rho](t)$ by solving (5), as well as the corresponding geometric density-to-current functional
$\mathbf{j}^{\rm GP}[\Psi_{0},\rho](t,\mathbf{r}):=\mathbf{j}_{\Psi^{\rm GP}[\Psi_{0},\rho](t)}(\mathbf{r}).$ (11)
It plays a central role in our new theory because, from the invertibility of $\mathcal{L}_{\Psi}$ in (9), we obtain that
$\partial_{t}\rho(t,\mathbf{r})+\nabla\cdot\mathbf{j}^{\rm GP}\big[\Psi_{0},\rho\big](t,\mathbf{r})=0\quad\Longleftrightarrow\quad\rho=\rho^{\rm S}.$ (12)
In other words, the sought-after $\rho^{\rm S}$ is the unique density that satisfies the standard continuity equation with the geometric current $\mathbf{j}^{\rm GP}[\Psi_{0},\rho](t,\mathbf{r})$. Eq. (12) is analogous to the van Leeuwen equation [29] in standard TDDFT. Note that the latter is second-order in time whereas (12) is first-order.
@@
In geometric Kohn–Sham theory, the exact density $\rho^{\rm S}$ is reproduced using $N$ non-interacting electrons described with a Slater determinant $\Phi:=(N!)^{-1/2}\det(\varphi_{n}(\mathbf{r}_{k}))$ evolving with an effective equation of the form $i\partial_{t}\Phi(t)=(\widehat{T}+\widehat{V}_{\rm ext}(t)+i\widehat{W}(t))\Phi(t)$ (conserving the normalization of $\Phi$), which can be re-written with a commutator as in (8)
$i\partial_{t}\varphi_{n}(t,\mathbf{r})=\Big(-\frac{\nabla^{2}}{2}+V_{\rm ext}(t,\mathbf{r})+i\big[W(t),\gamma_{\Phi(t)}\big]\Big)\varphi_{n}(t,\mathbf{r}),$ (13)
where $\gamma_{\Phi}(\mathbf{r},\mathbf{r}^{\prime}):=\sum_{n=1}^{N}\varphi_{n}(\mathbf{r})\varphi_{n}(\mathbf{r}^{\prime})^{\dagger}$ is the one-particle density matrix. In the sequel we will introduce several functionals $W$ to be used in (13) that all coincide for the solution reproducing the exact density $\rho^{\rm S}$.
@@
For an initial Slater determinant $\Phi_{0}$, we can apply the theory developed before with $\widehat{V}_{\rm ee}=0$ and obtain the non-interacting functionals $W_{0}[\Phi_{0},\rho]$ and $\mathbf{j}_{0}^{\rm GP}[\Phi_{0},\rho]$. We denote by $\Phi^{\rm GP}[\Phi_{0},\rho](t)$ the associated Slater determinant. If we pick $\rho=\rho^{\rm S}$, that is, solve (13) with $W=W_{0}[\Phi_{0},\rho^{\rm S}]$, where $\Phi_{0}$ has the same density as the correlated initial state $\Psi_{0}$, we obtain the trajectory $\Phi^{\rm GP}[\Phi_{0},\rho^{\rm S}](t)$ of Slater determinants reproducing the exact Schrödinger density $\rho^{\rm S}$. However, since $\rho^{\rm S}$ is unknown, we have to find a nonlinear equation involving only the Kohn–Sham state $\Phi$ or its density $\rho_{\Phi}$, giving $\Phi^{\rm GP}[\Phi_{0},\rho^{\rm S}](t)$ as unique solution. There are several possible such nonlinear equations, which we describe now.
Consider an arbitrary Slater trajectory $\Phi(t)$ starting with the density $\rho_{\Phi(0)}=\rho_{\Psi_{0}}$ and define
$\widetilde{W}[\Psi_{0},\Phi](t,\mathbf{r}):=\\ \frac{1}{2}\mathcal{L}_{\Phi(t)}^{-1}\nabla\cdot\big(\mathbf{j}_{\Phi(t)}(\mathbf{r})-\mathbf{j}^{\rm GP}[\Psi_{0},\rho_{\Phi}](t,\mathbf{r})\big),$ (14)
with $\mathcal{L}_{\Phi(t)}^{-1}$ the inverse of the operator in (10). Next insert this functional $\widetilde{W}[\Psi_{0},\Phi](t,\mathbf{r})$ in (13). We obtain a nonlinear equation in $\Phi$. Writing the continuity equation (9) for the solution $\widetilde{\Phi}(t)$ to this equation and using that $\widetilde{W}$ satisfies (14), we find that $\rho_{\widetilde{\Phi}(t)}$ solves Schrödinger’s continuity equation (12). This proves, as we wanted, that $\rho_{\widetilde{\Phi}}=\rho^{\rm S}$. In particular, $\widetilde{W}[\Psi_{0},\widetilde{\Phi}]=W_{0}[\Phi_{0},\rho^{\rm S}]$ and $\widetilde{\Phi}(t)=\Phi^{\rm GP}[\Phi_{0},\rho^{\rm S}](t)$ is the desired Slater solution. This is in the spirit of generalized Kohn–Sham theory [30], because we solve a nonlinear equation with $\widetilde{W}$ as a functional of $\Phi$ and not only the density $\rho_{\Phi}$.
To formulate a theory relying only on the density, we employ a manipulation similar to what is done for the study of the ground state, where the Hartree-exchange-correlation (Hxc) functional is defined as the difference between the interacting and non-interacting energies [31]. Inspired by [32, Sec. 1.3], we define the Geometric Kohn–Sham (GKS) functional by
$W^{\rm GKS}[\Psi_{0},\Phi_{0},\rho]:=W_{0}[\Phi_{0},\rho]-W[\Psi_{0},\rho].$ (15)
@@
This difference could also be expressed using the adiabatic connection formalism [33]. We then solve the resulting nonlinear equation (13) with $W=W^{\rm GKS}[\Psi_{0},\Phi_{0},\rho_{\Phi}]$. We call $\Phi^{\rm GKS}(t)$ the corresponding solution and $\rho^{\rm GKS}:=\rho_{\Phi^{\rm GKS}}$ its density. By uniqueness of $W$, we know that $W^{\rm GKS}[\Psi_{0},\Phi_{0},\rho^{\rm GKS}]$ must be the $W$ reproducing $\rho^{\rm GKS}$, which is by definition $W_{0}[\Phi_{0},\rho^{\rm GKS}]$. But then from (15) we obtain $W[\Psi_{0},\rho^{\rm GKS}]=0$ and therefore $\rho^{\rm GKS}=\rho^{\rm S}$ and $\Phi^{\rm GKS}=\Phi^{\rm GP}[\Phi_{0},\rho^{\rm S}]$, as we wanted.
We have seen that the Slater trajectory $\Phi^{\rm GP}[\Phi_{0},\rho^{\rm S}](t)$ reproducing the exact density $\rho^{\rm S}$ solves two different equations with a nonlinear term involving either $\Phi$ or $\rho_{\Phi}$. There are other choices. For instance, one can also consider the functional defined for a given $\rho(t,\mathbf{r})$ by
$\widetilde{W}^{\rm GKS}[\Psi_{0},\Phi_{0},\rho](t,\mathbf{r}):=\\ \frac{1}{2}\mathcal{L}_{\Phi^{\rm GP}[\Phi_{0},\rho](t)}^{-1}\nabla\!\cdot\!\Big(\mathbf{j}^{\rm GP}_{0}[\Phi_{0},\rho](t,\mathbf{r})-\mathbf{j}^{\rm GP}[\Psi_{0},\rho](t,\mathbf{r})\Big),$ (16)
obtained by replacing $\Phi(t)$ by $\Phi^{\rm GP}[\Phi_{0},\rho](t)$ in (14). This involves the difference of the currents instead of that of the $W$’s. The choice of a given functional should be motivated by how easily approximations can be made. This is not the focus of this Letter.

### v2

@@ Abstract
We provide a new formulation of Time-Dependent Density Functional Theory (TDDFT) based on the geometric structure of the set of states constrained to have a fixed density. Orbital-free TDDFT is formulated using a hydrodynamics equation involving a new density-to-current functional map. In the corresponding Kohn–Sham equation, the density is reproduced using a non-local operator. Finally, we present numerical simulations for one-dimensional soft-Coulomb systems.
Density Functional Theory (DFT) is very successful in describing electrons at equilibrium, that is, close to their ground states. It is nowadays the reference computational method for large molecules and solids, both in academic research and industrial applications. By contrast, Time-Dependent Density Functional Theory (TDDFT) [1, 2, 3] is less developed and still suffers from important limitations [4, 5, 6]. It is however the only viable computational approach for complex systems at present, and the demand for novel concepts and approximations in TDDFT is therefore considerable. Practical applications of TDDFT include charge transfer in complex molecules [7], electronic response of large systems [8], and ultrafast phenomena in attosecond physics [9]. In this Letter we open a new avenue in TDDFT. Based on geometric concepts, we introduce a new exact scheme that seems to better describe systems driven out of equilibrium. We also present 1D numerical simulations that illustrate the potential power of our new approach, when used as a correction to adiabatic density functionals.
@@
where $\widehat{T}:=-\sum_{j=1}^{N}\nabla^{2}_{\mathbf{r}_{j}}/2$ is the kinetic-energy operator, $\widehat{V}_{\rm ee}$ is the electron-electron interaction and $\widehat{V}_{\rm ext}(t):=\sum_{j=1}^{N}V_{\rm ext}(t,\mathbf{r}_{j})$ is a time-dependent external potential. The reference $N$-particle Schrödinger equation is
@@
In this first section, we focus on orbital-free TDDFT, where the high-dimensional linear equation (2) is replaced by an equivalent nonlinear equation involving only the one-particle density
@@
To this end, we consider modified Schrödinger equations of the form
$i\partial_{t}\Psi(t)=\Big(\widehat{H}(t)+\widehat{F}[\Psi_{0},V_{\rm ext},\rho](t)\Big)\Psi(t),\quad\Psi(0)=\Psi_{0},$ (4)
where the correction term $\widehat{F}[\Psi_{0},V_{\rm ext},\rho](t)$ is such that the density $\rho_{\Psi(t)}(\mathbf{r})$ of the solution is equal to a density $\rho(t,\mathbf{r})$ that is prescribed in advance. The correction term is a functional of the density $\rho$, of the time-dependent external potential $V_{\rm ext}$ and of the initial state $\Psi_{0}$ which is of course assumed to have density $\rho(0,\mathbf{r})$. To ensure causality, $\widehat{F}[\Psi_{0},V_{\rm ext},\rho](t)$ is assumed to only depend on $\rho(s,\mathbf{r})$ and $V_{\rm ext}(s,\mathbf{r})$ for past times $0\leqslant s\leqslant t$.
@@
There are many possible choices for $\widehat{F}$. Based on geometric considerations, we will construct in this section a functional for which the density $\rho^{\rm S}:=\rho_{\Psi^{\rm S}}$ of the Schrödinger wavefunction $\Psi^{\rm S}$ in (2) is the only solution to the implicit equation $\widehat{F}[\Psi_{0},V_{\rm ext},\rho]=0$. This is how (2) can be reformulated in terms of the density only. In the next section, we consider the corresponding Kohn–Sham theory which provides $\rho^{\rm S}$ in terms of orbitals obtained from a non-interacting Hamiltonian with a similar but different correction $\widehat{F}$ that is non-zero for $\rho=\rho^{\rm S}$.
@@
Standard TDDFT can be reinterpreted as the choice $\widehat{F}^{\rm VP}[\Psi_{0},V_{\rm ext},\rho](t):=\sum_{j=1}^{N}V(t,\mathbf{r}_{j})$ with a potential functional $V=V[\Psi_{0},V_{\rm ext},\rho]$. This can be achieved by requiring the action to be stationary with respect to wavefunction variations at fixed density [10, Sec. III.A], which we refer to as the Variational Principle (VP). The Runge–Gross theorem [1] states that $V$ is unique modulo a time-dependent constant, hence $\rho^{\rm S}$ is the unique solution to $V[\Psi_{0},V_{\rm ext},\rho]=0$. Since $V[\Psi_{0},V_{\rm ext},\rho]=V[\Psi_{0},0,\rho]-V_{\rm ext}$, this last condition can be formulated with the universal functional $V[\Psi_{0},\rho]:=V[\Psi_{0},0,\rho]$ as
$V[\Psi_{0},\rho](t,\mathbf{r})=V_{\rm ext}(t,\mathbf{r})+C(t)\quad\Longleftrightarrow\quad\rho=\rho^{\rm S},$ (5)
where $C(t)$ is a time-dependent constant.
@@
In our new Geometric Principle (GP) we do not demand the action to be stationary, but instead require that the norm of the correction, $\|\widehat{F}^{\rm GP}[\Psi_{0},V_{\rm ext},\rho]\Psi(t)\|=\|i\partial_{t}\Psi(t)-\widehat{H}(t)\Psi(t)\|$, be the smallest it can be at all times. For simplicity, we first consider a prescribed time-independent density $\rho(\mathbf{r})$. The wavefunction $\Psi(t)$ in (4) then evolves on the set $\mathcal{M}_{\rho}$ of all wavefunctions having this prescribed $\rho$ and the velocity $\partial_{t}\Psi(t)$ must belong to the tangent space $\mathcal{T}_{\Psi(t)}\mathcal{M}_{\rho}$ at all times. As shown in Fig. 1, the correction term has minimal norm when $\partial_{t}\Psi(t)$ is the projection of $i\widehat{H}(t)\Psi(t)$ onto the tangent space, that is, when $i\widehat{F}^{\rm GP}[\Psi_{0},V_{\rm ext},\rho]\Psi(t)$ is in the normal space $\mathcal{N}_{\Psi(t)}:=(\mathcal{T}_{\Psi(t)}\mathcal{M}_{\rho})^{\perp}$. By simple calculations [10, Sec. II.B], the normal space is
@@
and the new correction term is thus
$\widehat{F}^{\rm GP}[\Psi_{0},V_{\rm ext},\rho]\Psi(t):=i\sum_{j=1}^{N}W(t,\mathbf{r}_{j})\Psi(t),$ (7)
with a real-valued function $W(t,\mathbf{r})$. For a prescribed time-dependent density, the displacement of the tangent space in the normal direction only depends on $\partial_{t}\rho(t)$ and does not change the form of the correction in (7[10, Sec. II.C]. By construction, $W(t,\mathbf{r})$ only depends on $\Psi(t)$, $V_{\rm ext}(t)$ and $\partial_{t}\rho(t)$ and is thus causal. Hence the modified Schrödinger equation (4) is
$i\partial_{t}\Psi(t)=\left(\widehat{H}(t)+i\widehat{W}(t)\right)\Psi(t),$ (8)
where $\widehat{W}(t):=\sum_{j=1}^{N}W(t,\mathbf{r}_{j})$ and $W=W[\Psi_{0},V_{\rm ext},\rho]$ is a real-valued functional.
Equation (8) looks like a Schrödinger equation with a complex potential, similar to those considered in TDDFT for describing resonances [11, 12, 13, 14, 15] or open systems [16, 17, 18]. However, here, $W$ is chosen so that $\rho_{\Psi(t)}=\rho(t)$ and thus the resulting $\Psi(t)$ stays normalized for all times. This means that $\langle\Psi(t),\widehat{W}(t)\Psi(t)\rangle=0$, which implies that Eq. (8) can be rewritten in commutator form
$i\partial_{t}\Psi(t)=\bigg(\widehat{H}(t)+i\Big[\widehat{W}(t)\,,\,|\Psi(t)\rangle\langle\Psi(t)|\Big]\bigg)\Psi(t).$ (9)
For any fixed $\widehat{W}(t)$ and $\Psi(t)$, the modified Hamiltonian in (9) is Hermitian, and thus manifestly norm preserving.
The idea of constraining a Schrödinger dynamics using a projection on the tangent space, often attributed to McLachlan [19], can be traced back to Dirac [20] and Frenkel [21]. Depending on the considered constraints, the GP can lead to the same solution as the VP, or to completely different solutions [22, 23, 24, 25, 26, 27]. For a fixed density the GP and VP corrections and wavefunctions are very different [10, Secs. IV.A, VIII.C-D].
In [10, Thm. 3] we give a rigorous proof of the existence and uniqueness of $W[\Psi_{0},V_{\rm ext},\rho]$ in finite dimension. In Appendix A we rigorously prove a Runge–Gross theorem for Eq. (8) with smooth potentials. Hence $W[\Psi_{0},V_{\rm ext},\rho]$ is uniquely defined, when it exists, and $\rho^{\rm S}$ is the unique solution to $\widehat{F}^{\rm GP}[\Psi_{0},V_{\rm ext},\rho]=0$, that is,
$W[\Psi_{0},V_{\rm ext},\rho]=0\quad\Longleftrightarrow\quad\rho=\rho^{\rm S}.$ (10)
The function $W$ is very different in nature from the potential $V$ used in standard TDDFT. In particular, the potential $V$ often contains high peaks and steps to force the particles to move to the desired region of space [28]. This only works thanks to a complicated interplay between $V$ and the kinetic-energy operator $\widehat{T}$. Indeed, if we remove $\widehat{T}$, then adding a real-valued local potential $V$ modifies the phase of the wavefunction but not the density. In constrast, the function $W$ appears directly in the continuity equation associated with (8) or (9),
@@
where $\mathbf{j}_{\Psi}$ and $\rho_{\Psi}^{(2)}$ are the current and the two-particle density of $\Psi$. The linear operator $\mathcal{L}_{\Psi}$ is expected to be invertible in all cases of physical interest [10, Sec. VII.B]. The interpretation of (11) is that $W$ gives rise to sources and sinks [12], which is a more direct and flexible way of changing the density. In fact, on a lattice, densities moving too fast are not $V$-representable due to the finite speed of propagation [13] but can perfectly be $W$-representable [10, Secs. VII.C and VII.D].
Let us now re-interpret Eq. (10) in light of the continuity equation (11). From the functional $W[\Psi_{0},V_{\rm ext},\rho]$ we obtain a unique wavefunction $\Psi^{\rm GP}[\Psi_{0},V_{\rm ext},\rho](t)$ by solving (8) or (9), as well as the corresponding geometric density-to-current functional
$\mathbf{j}^{\rm GP}[\Psi_{0},V_{\rm ext},\rho](t,\mathbf{r}):=\mathbf{j}_{\Psi^{\rm GP}[\Psi_{0},V_{\rm ext},\rho](t)}(\mathbf{r}),$ (13)
the current of the wavefunction $\Psi^{\rm GP}[\Psi_{0},V_{\rm ext},\rho](t)$. It plays a central role in our new theory because, from the invertibility of $\mathcal{L}_{\Psi}$ in (11), we can rewrite Eq. (10) as
$\partial_{t}\rho(t,\mathbf{r})+\nabla\cdot\mathbf{j}^{\rm GP}\big[\Psi_{0},V_{\rm ext},\rho\big](t,\mathbf{r})=0\quad\Longleftrightarrow\quad\rho=\rho^{\rm S}.$ (14)
In other words, the sought-after $\rho^{\rm S}$ is the unique density that satisfies the standard continuity equation with the geometric current $\mathbf{j}^{\rm GP}[\Psi_{0},V_{\rm ext},\rho](t,\mathbf{r})$. For $\rho=\rho^{\rm S}$, we recover the Schrödinger wavefunction and current, $\Psi^{\rm GP}[\Psi_{0},V_{\rm ext},\rho^{\rm S}]=\Psi^{\rm S}$ and $\mathbf{j}^{\rm GP}\big[\Psi_{0},V_{\rm ext},\rho^{\rm S}\big]=\mathbf{j}_{\Psi^{\rm S}(t)}$. Equation (14) is our new geometric formulation of orbital free TDDFT. It is analogous to the van Leeuwen equation [29] in standard TDDFT. Note that the latter is second-order in time whereas (14) is first-order.
@@
In geometric Kohn–Sham theory, the exact density $\rho^{\rm S}$ is reproduced using $N$ non-interacting electrons described with a Slater determinant $\Phi(t):=(N!)^{-1/2}\det(\varphi_{n}(t,\mathbf{r}_{k}))$ with the orbitals $\varphi_{n}$ solving the equation
$i\partial_{t}\varphi_{n}(t,\mathbf{r})\!=\!\Big(-\frac{\nabla^{2}}{2}+V_{\rm ext}(t,\mathbf{r})+i\big[W^{\rm KS}(t),\gamma_{\Phi(t)}\big]\Big)\varphi_{n}(t,\mathbf{r}),$ (15)
where $\gamma_{\Phi(t)}(\mathbf{r},\mathbf{r}^{\prime}):=\sum_{n=1}^{N}\varphi_{n}(t,\mathbf{r})\varphi_{n}(t,\mathbf{r}^{\prime})^{\dagger}$ is the one-particle density matrix. The correction term is of commutator form as in (9) with $W^{\rm KS}$ a functional to be specified so that the exact Schrödinger density is reproduced, i.e. $\rho_{\Phi(t)}=\rho^{\rm S}(t)$. To obtain a predictive scheme, this functional should of course not depend on the unknown density $\rho^{\rm S}$.
@@
For an initial Slater determinant $\Phi_{0}$, we can apply the theory developed before with $\widehat{V}_{\rm ee}=0$ and obtain the corresponding non-interacting functional $W_{0}[\Phi_{0},V_{\rm ext},\rho]$. If we pick $\rho=\rho^{\rm S}$, that is, solve (15) with $W^{\rm KS}=W_{0}[\Phi_{0},V_{\rm ext},\rho^{\rm S}]$, where $\Phi_{0}$ has the same density as the initial state $\Psi_{0}$, we obtain the trajectory of Slater determinants reproducing the exact Schrödinger density $\rho^{\rm S}$. In practice we do not know $\rho^{\rm S}$ and we can only conclude that the sought-after functional $W^{\rm KS}$ in Eq. (15) must coincide with $W_{0}[\Phi_{0},V_{\rm ext},\rho^{\rm S}]$ for (and only for) $\rho_{\Phi}=\rho^{\rm S}$. There are several choices for $W^{\rm KS}$.
Recall that in the Kohn–Sham scheme of standard TDDFT, the universal Hartree-exchange-correlation (Hxc) potential functional is defined as [31, Sec. 1.3]
$V_{\rm Hxc}[\Psi_{0},\Phi_{0},\rho]:=V_{0}[\Phi_{0},\rho]-V[\Psi_{0},\rho],$ (16)
where $V_{0}[\Phi_{0},\rho]$ is the universal Kohn–Sham potential and $V[\Psi_{0},\rho]$ the universal interacting potential functional as above. In the same spirit, we define here
$\widetilde{W}^{\rm KS}[\Psi_{0},\Phi_{0},V_{\rm ext},\rho]:=W_{0}[\Phi_{0},V_{\rm ext},\rho]-W[\Psi_{0},V_{\rm ext},\rho].$ (17)
@@
Let us call $\Phi(t)$ the solution to the nonlinear equation (15) with $W^{\rm KS}=\widetilde{W}^{\rm KS}[\Psi_{0},\Phi_{0},V_{\rm ext},\rho_{\Phi}]$. By uniqueness in Thm. 1 (b) below, we know that this $W^{\rm KS}$ must be the one reproducing $\rho_{\Phi}$, which is by definition $W_{0}[\Phi_{0},V_{\rm ext},\rho_{\Phi}]$. But then from (17) we obtain $W[\Psi_{0},V_{\rm ext},\rho_{\Phi}]=0$ and therefore $\rho_{\Phi}=\rho^{\rm S}$ as we wanted.
The functional introduced in Eq. (17) is a natural extension of Eq. (16) but it implicitly depends on the external potential $V_{\rm ext}$. To turn it into a universal functional we can replace $V_{\rm ext}$ by $V[\Psi_{0},\rho]$ since for $\rho^{\rm S}$ we will have equality by (5). By definition any density $\rho$ is exactly reproduced with $V[\Psi_{0},\rho]$, hence no $W$ is necessary and we have $W[\Psi_{0},V[\Psi_{0},\rho],\rho]=0$. Thus, Eq. (17) leads to the universal geometric Kohn–Sham functional
$W^{\rm KS}[\Psi_{0},\Phi_{0},\rho]:=W_{0}\big[\Phi_{0},V[\Psi_{0},\rho],\rho\big],$ (18)
that is now independent of $V_{\rm ext}$. The solution $\Phi(t)$ to the nonlinear equation (15) with $W^{\rm KS}=W^{\rm KS}[\Psi_{0},\Phi_{0},\rho_{\Phi}]$ has the desired density $\rho^{\rm S}$. Indeed, by uniqueness in Thm. 1 (b) we know that this $W^{\rm KS}$ must be the one reproducing $\rho_{\Phi}$ in the external potential $V_{\rm ext}$, that is, $W_{0}\big[\Phi_{0},V[\Psi_{0},\rho_{\Phi}],\rho_{\Phi}\big]=W_{0}[\Phi_{0},V_{\rm ext},\rho_{\Phi}]$. But this equality means that the two potentials $V[\Psi_{0},\rho_{\Phi}]$ and $V_{\rm ext}$ yield the same $\rho_{\Phi}$ in presence of this $W^{\rm KS}$. By uniqueness of $V$ at given $W$ as stated in Thm. 1 (a) below, we conclude that $V[\Psi_{0},\rho_{\Phi}]=V_{\rm ext}+C(t)$. By (5) we finally obtain $\rho_{\Phi}=\rho^{\rm S}$ as desired.
We have introduced two nonlinear functionals to be used in the geometric Kohn–Sham scheme (15), that reproduce the exact Schrödinger density $\rho^{\rm S}$. There are other possibilities. The choice of a given functional should be motivated by how easily approximations can be made. This is not the focus of this Letter.